OpenStudy (anonymous):

Proove the double angle formula ''sin 2x = 2sin x cosx '' from the addition formula : sin(x+y) = sin x cos y+ cos x sin y

OpenStudy (anonymous):

No idea, I do know how to do it from a different formula.

myininaya (myininaya):

sin(x+x)=sinxcosx+cosxsinx=2sinxcosx

myininaya (myininaya):

which =sin(2x)

myininaya (myininaya):

since x+x=2x

myininaya (myininaya):

this is all?

OpenStudy (anonymous):

Tried using euler's formula?

OpenStudy (anonymous):

eum, well, i think it is right, but I think you've made a mistake in the beginging, it is sin(x+y) not sin(x+x)

OpenStudy (anonymous):

I don't know the euler's formula

OpenStudy (anonymous):

\[e^{ix} = \cos{x}+i\sin{x}\]

myininaya (myininaya):

but you want to use the addition rule to prove sin(2x)=2sin(x)cos(x) sin(2x)=sin(x+x)=sin(x)cos(x)+sin(x)cos(x)=2sin(x)cos(x)

myininaya (myininaya):

this is all you have to do just use that formula just like it says

OpenStudy (anonymous):

Sorry, didn't read the question properly, myininaya's looks good to me.